Determine Two Integers Whose Sum Is 4 And Difference Is 4.
Found 2 solutions by JulietG, richwmiller:
Answer by JulietG(1812)

Here is the set of possible pairs that equal 4:
0,4
1,3
2,2
Which of those has a difference of 4? Only one of them ;-)

Answer by richwmiller(17219)

Another way
a+b=4
a-b=4
add them
2a=8
a=4
4+b=4
b=0